# Enhancing Fault Detection for Large Language Models via Mutation-Based Confidence Smoothing

### v1

@@ 6.3. RQ3: Effectiveness of MuCS
Table 4 and Table 5 summarize the TRC scores of various methods enhanced by MuCS (with -M as a suffix), and the relative difference between the scores with and without using MuCS. The results demonstrate that the MuCS significantly boosts the performance of fault detection. In terms of the average TRC scores of each method on different tasks, in 19 out of 20 cases, the performance of existing fault detection methods has been increased with TRC improvements ranging from 0.0366 to 0.1794. Considering the average relative improvement, confidence smooth can enhance the performance of existing methods by up to 97.64% (LLaMA for sentiment analysis, method: Margin). Therefore, we can conclude that MuCS is effective in boosting the existing fault detection methods on LLMs.
Comparing the improvements across different LLMs, we found that improvements in GPT4 are weaker than improvements in other LLMs. On average, the improvements in LLaMA, GPT3.5, and GPT4 are 37.63%, 41.41%, and 11.78%, respectively. We conjecture that the reason for this phenomenon is that GPT4 is more powerful and more resilient to the input mutation. Thus, the smooth is less useful for GPT4. To check our assumption, we compute the average variance of the prediction output probability changes after prompt mutation of each LLM. The results show that the probability changes of GPT are negligible. For example, for the sentiment analysis task, the variance of the probability changes of LLaMA, GPT3.5, and GPT4 are 0.0047, 0.0158, and 0.0041, respectively.
Finally, even though the mutation-based confidence smooth can enhance the fault detection of LLMs, the best methods only have 0.5472, 0.6769, and 0.6687 TRC scores in LLaMA, GPT3.5, and GPT4, respectively. There is still a big room between the TRC scores achieved by existing methods and the ideal performance.
Answer to RQ3: MuCS significantly enhances the effectiveness of existing fault detection methods by up to 97.62% in terms of the TRC score.
@@ 6.4. RQ4: MuCS Affected Prediction Confidence
Model | Confidence () | ECE () | Diversity () | |||
Before | After | Before | After | Before | After | |
CD-GPT35 | 0.8535 | 0.8863 | 0.1915 | 0.2563 | 584.96 | 655.29 |
CD-GPT4 | 0.8575 | 0.8665 | 0.1630 | 0.1638 | 696.62 | 542.42 |
PC-GPT35 | 0.7659 | 0.7349 | 0.4560 | 0.3449 | 432.22 | 40.89 |
TMN-GPT35 | 0.6013 | 0.5981 | 0.0982 | 0.1030 | 205.36 | 24.69 |
TMN-GPT4 | 0.8187 | 0.7571 | 0.0407 | 0.0846 | 196.56 | 62.29 |
Sentiment-LLaMA | 0.5476 | 0.5626 | 0.3065 | 0.3041 | 92.67 | 67.00 |
Sentiment-GPT35 | 0.8596 | 0.8153 | 0.0917 | 0.1180 | 335.87 | 104.87 |
Sentiment-GP4 | 0.8320 | 0.8059 | 0.0787 | 0.1093 | 230.60 | 93.33 |
Average change | 1.78% | 4.04% | 42.67% |
We further study how MuCS affects the prediction of LLMs and whether it diversifies the prediction confidence or not. Figure 5 depicts the confidence of LLMs on the datasets before and after confidence smoothing and Table 6 summarizes the average confidence and ECE scores. The results show no big difference between the average confidence of LLMs and ECE scores before and after smoothing. This finding is consistent with the conclusion from a previous work (Zhu et al., 2023b) that states the effectiveness of fault detection methods is not directly related to the calibration of the model (most confidence calibration methods are useless or harmful for fault detection).
Then, we check whether MuCS can increase the diversity of confidence distribution or not. We compute the variance of the number of confidence scores in each range to check this diversity. The results presented in Table 6 demonstrated that, except for the clone detection tasks, LLMs have more diverse predictions on the datasets with an average 42.67% diversity improvement. Taking Figure 6(b) as an example, the confidence scores of GPT3.5 are concentrated in 1.0 which is highly contradictory to its accuracy (36%). After the mutation-based smooth, confidence scores are spread throughout the distribution which is more useful to help understand the capability of LLMs on this task.
Answer to RQ4: MuCS significantly increases the diversity of confidence distribution of LLMs with an average 42.67% improvement and eliminates the concentrated confidence problem.
@@ 7.1. Limitation and Future Direction
Limitation. The potential limitations are that 1) we only consider classification tasks. Non-classification tasks such as question answering are not included in this work. The reason is that most of the existing fault detection methods are designed for classification tasks only. Even though there also exist methods that can be employed for non-classification tasks, such as PRIMA (Wang et al., 2021), applying them for LLMs is not practical. PRIMA needs to mutate deep learning models, however, as LLMs are much larger than classical deep learning models, it is difficult to mutate LLMs and save the mutants. As a result, only classification tasks are studied in our work. 2) For a given input, our proposed mutation-based confidence smooth solution needs to access LLMs times, thus, the money consumption is times than the normal way when we test closed-source LLMs.
Future direction. As LLMs (such as ChatGPT) become the most used deep learning models nowadays, the most important task people would use is the generation tasks, e.g., chatting using ChatGPT. One of the most important directions in this field is to propose fault detection methods for non-classification tasks, especially generation tasks, and then use these methods to help reveal faults in LLMs. Some metrics (Kuhn et al., 2023; Huang et al., 2023) have been designed for measuring the uncertainty in LLMs, how to combine these uncertainty metrics with existing methods could be the first try.
@@ 7.2. Threats to Validity
The internal threat lies in the implementation of fault detection methods, LLMs, and prompt mutation methods. The implementation of each fault detection method comes from the official project provided original paper. The implementation of LLMs comes from the famous open-source project Hugging Face. The code of text mutation also comes from the commonly used project TextAugment (Marivate and Sefara, 2020). The implementation of code refactoring methods modified from the previous work (Dong et al., 2023b).
The external threat comes from our studied tasks, datasets, LLMs, and fault detection methods. As mentioned in Section 7.1, we only consider the classification tasks in this work. For the tasks, we consider tasks from both NLP and SE fields. Two traditional NLP tasks and two programming tasks are included in our work. For the studied LLMs, we consider both open-source LLMs, e.g., LLaMA, and closed-source LLMs (GPT4), where GPT4 is recognized as the SOTA LLM currently. Besides, another threat that comes from the datasets and models is that there might be data leakage issues in LLMs. For example, we found GPT4 has 100% accuracy on the problem classification task which means our test set is likely in the training set of GPT4. However, it is difficult to check this issue as GPT3.5 and GPT4 are closed-source LLMs. We can only exclude the cases from the analysis of the accuracy of LLMs. For the fault detection methods, we follow the previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) and explore 10 fault detection methods including both learning-based methods and output-based methods in our study. These methods are from different communities, for example, DeepGini is from the SE community and TestRank is from the ML community. We believe our studied objectives are representative and diverse enough and the conclusion drawn from this work can be generalized to other similar cases.
The construct threat could be the hyperparameter settings and the non-determinism of LLMs. We directly use the default settings suggested by Hugging Face and OpenAI’s API to build or access LLMs. Besides, we release the datasets and prompts used in our experiments to support to reproduce results reported in our work and for future research.
@@ 8. Conclusion
In this work, we first explored the potential of fault detection methods for LLMs. Based on the experiments on four tasks including NLP and code tasks, four LLMs, and nine fault detection methods, we found LLMs are overconfident with their predictions and current fault detection methods are not effective in revealing faults in LLMs. To enhance the detection performance, we proposed MuCS, which uses mutation analysis to smooth the prediction confidence of LLMs. Specifically, we utilized text transformation and code refactoring techniques to mutate the inputs and collect the averaged output probabilities of all mutants as the final prediction confidence of LLMs. Evaluation results demonstrated that MuCS significantly increased the diversity of output distribution of LLMs and enhanced the effectiveness of fault detection by up to 97.64% in terms of test relative coverage score.

### v2

@@ 6.3. RQ3: Effectiveness of MuCS
Table 3 summarizes the average TRC scores achieved by different fault detection methods across all considered labeling budgets using MuCS (with -M as a suffix), and the relative difference between the scores with and without using MuCS. The results clearly demonstrate that the MuCS significantly boosts the performance of fault detection. Specifically, in 52 out of 71 cases, the performance of existing fault detection methods has been increased with TRC improvements ranging from 0.48% to 70.58%. Considering the average results (the last row in the Table), MuCS can enhance all fault detection methods with improvement by up to 13.57%. Besides, we conducted the statistical analysis using two methods Welch’s t-test and Wilcoxon signed-rank test to investigate the significance of improvements introduced by MuCS. Concretely, we collect all TRC scores resulting from each method with and without using MuCS and compute the significance. The results are shown in Table 4. The t-test suggests that for four methods (Gini, MCP, MaxP, and ATS), the improvements are significant, and the signed-rank test demonstrates all improvements are significant (with P-value ¡ 0.03). Therefore, we can conclude that MuCS is effective in boosting the existing fault detection methods on LLMs.
Comparing the improvements across different LLMs, we found that improvements in closed-source LLMs are greater than improvements in open-source LLMs. On average, the improvements in DeepSeekCoder, LLaMA3, GPT3.5, and GPT4 are 3.82%, 3.22%, 16.38%, and 7.87%, respectively. This means MuCS is an effective framework to enhance the black-box testing of LLMs. Then, after comparing each method after using MuCS, we found that MaxP, another simple method that directly utilizes the top-output probability as the indicator for data prioritization performs the best. Combining the conclusion in Section 4.2, simple methods perform better than learning-based methods with and without using MuCS.
Finally, even though the mutation-based confidence smooth can enhance the fault detection for LLMs, the best method only has a 0.7641 average TRC score. There is still a big room between the TRC scores achieved by existing methods and the ideal performance.
Answer to RQ3: MuCS significantly enhances the effectiveness of existing fault detection methods by up to 70.53% in terms of the TRC score.
@@ 6.4. RQ4: MuCS Affected Prediction Confidence
| ECE () | Diversity () | |||
| Model | Before | After | Before | After |
| CD-DeepSeekCoder | 0.9421 | 0.9112 | 11066.62 | 9535.22 |
| CD-GPT35 | 0.1915 | 0.2563 | 584.96 | 655.29 |
| CD-GPT4 | 0.1630 | 0.1638 | 696.62 | 542.42 |
| PC-DeepSeekCoder | 0.3110 | 0.1713 | 7963.02 | 1189.08 |
| PC-GPT35 | 0.4560 | 0.3449 | 205.36 | 24.69 |
| PC-GPT4 | 0.0143 | - | 332.00 | - |
| TMN-LLaMA3 | 0.1711 | 0.0701 | 12338.18 | 1641.37 |
| TMN-GPT35 | 0.0982 | 0.1030 | 205.36 | 24.69 |
| TMN-GPT4 | 0.0407 | 0.0846 | 196.56 | 62.29 |
| Sentiment-LLaMA3 | 0.5714 | 0.2345 | 8301.62 | 4898.62 |
| Sentiment-GPT35 | 0.0917 | 0.1180 | 335.87 | 104.87 |
| Sentiment-GP4 | 0.0787 | 0.1093 | 230.60 | 93.33 |
| Average change | 17.60% | 55.44% | ||
We further study how MuCS affects the prediction of LLMs and whether it diversifies the prediction confidence or not. Figure 5 depicts the confidence of LLMs on the datasets before and after confidence smoothing and Table 5 summarizes the average ECE scores. The results show that even though MuCS can reduce the average ECE scores (i.e., better calibration), there are still multiple pairs that have similar ECE scores, e.g., Sentiment-GPT35 and Sentiment-GPT4. Moreover, previous works (Zhu et al., 2023b) stated the effectiveness of fault detection methods is not directly related to the calibration of the model (most confidence calibration methods are useless or harmful for fault detection). It is not clear whether the improvements are introduced by better calibration.
Then, we check if MuCS can increase the diversity of confidence distribution to solve the confidence concentrated issue. We compute the variance of the number of confidence scores in each range to check this diversity. The results presented in Table 5 demonstrated that, except for one case CD-GPT35, LLMs have more diverse predictions on the datasets with an average 55.44% diversity improvement. Taking Figure 5(f) as an example, the confidence scores of GPT3.5 are concentrated in 1.0 which is highly contradictory to its accuracy (36%). After the mutation-based smooth, confidence scores are spread throughout the distribution which is more useful to help understand the capability of LLMs for this task.
Answer to RQ4: MuCS significantly increases the diversity of confidence distribution of LLMs with an average 55.44% improvement and eliminates the concentrated confidence problem.
@@ 7. Discussion
7.1. Confidence vs. Correctness
| Without MuCS | With MuCS | |||
| Model | Significance | P-value | Significance | P-value |
| CD-DeepSeekCoder | 0.2753 | 7.43E-19 | 0.2243 | 1.09E-12 |
| CD-GPT3.5 | 0.2636 | 1.21E-13 | 0.3032 | 9.41E-18 |
| CD-GPT4 | 0.2969 | 4.71E-17 | 0.2823 | 1.70E-15 |
| PC-DeepSeekCoder | -0.0522 | 9.92E-02 | 0.2402 | 1.38E-14 |
| PC-GPT3.5 | 0.0952 | 5.22E-02 | 0.2546 | 1.07E-07 |
| TMN-LLaMA3 | 0.1110 | 2.53E-05 | 0.2426 | 1.21E-20 |
| TMN-GPT3.5 | 0.2904 | 2.97E-29 | 0.3238 | 2.47E-36 |
| TMN-GPT4 | 0.3909 | 1.55E-53 | 0.4059 | 5.97E-58 |
| Sentiment-LLaMA3 | 0.4603 | 1.34E-53 | 0.3730 | 2.33E-34 |
| Sentiment-GPT3.5 | 0.2007 | 4.37E-03 | 0.2124 | 2.53E-03 |
| Sentiment-GPT4 | 0.4263 | 3.10E-10 | 0.4332 | 1.49E-10 |
As mentioned before, all considered fault detection methods rely on the assumption – there is a connection between the prediction confidence and the prediction correctness. However, it is unclear whether the assumption stands or not, especially for LLMs. We design an exploratory study to answer this problem. Specifically, we label the corrected prediction as 1 and the incorrect prediction as 0 and then utilize Pearson correlation coefficient to quantify the relationship between confidence and correctness as shown in Table 6. We can see that there is a clear positive correlation between these two indicators except for the problem classification task before using MuCS. Moreover, after employing MuCS, all correlations become positive with a significance less than 0.03. This exploratory study demonstrates the relationship between confidence and correctness exists, and can be enhanced by MuCS.
@@ 7.2. Limitation and Future Direction
Limitation. For a given input, our proposed mutation-based confidence smooth solution needs to access LLMs times, thus, the money consumption is times than the normal way when we test closed-source LLMs. However, from the view of executing time, the increased cost is negligible since it is possible to perform the prediction of a batch of data at the same time. There is indeed a trade-off between the monetary cost of accessing close-source LLMs and data labeling. This should be considered case by case.
Future direction. In this work, we only consider classification tasks since most existing fault detection methods are proposed for these tasks. Some methods, such as PRIMA (Wang et al., 2021) can be used for none classification tasks directly. Unfortunately, applying it to LLMs is not practical as it needs to mutate deep learning models but LLMs are much larger than classical deep learning models. However, the most important tasks people would use are generation tasks, e.g., chatting using ChatGPT. One of the important directions in this field is to propose fault detection methods for non-classification tasks, especially generation tasks, and then use these methods to help reveal critical faults in LLMs. Some metrics (Kuhn et al., 2023; Huang et al., 2023) have been designed for measuring the uncertainty in LLMs, how to combine these uncertainty metrics with existing methods could be the first try.
@@ 7.3. Threats to Validity
The internal threat lies in the implementation of fault detection methods, LLMs, and prompt mutation methods. The implementation of each fault detection method comes from the official project provided in the original paper. The implementation of LLMs comes from the famous open-source project Hugging Face. The code of text mutation also comes from the commonly used project TextAugment (Marivate and Sefara, 2020). The implementation of code refactoring methods modified from the previous work (Dong et al., 2023b).
The external threat comes from our studied tasks, datasets, LLMs, and fault detection methods. As mentioned in Section 7.2, we only consider the classification tasks in this work. For the tasks, we consider tasks from both NLP and SE fields. Two traditional NLP tasks and two programming tasks are included in our work. For the studied LLMs, we consider both open-source LLMs, e.g., LLaMA, and closed-source LLMs (GPT4), where GPT4 is recognized as the SOTA LLM currently. Besides, another threat that comes from the datasets and models is that there might be data leakage issues in LLMs. For example, we found GPT4 has 100% accuracy on the problem classification task which means our test set is likely in the training set of GPT4. However, it is difficult to check this issue as GPT3.5 and GPT4 are closed-source LLMs. We can only exclude the cases from the accuracy analysis of LLMs. For the fault detection methods, we follow the previous works (Hu et al., 2024; Ma et al., 2021; Hu et al., 2023) and explore 10 fault detection methods including both learning-based methods and output-based methods in our study. These methods are from different communities, for example, DeepGini is from the SE community and TestRank is from the ML community. We believe our studied objectives are representative and diverse enough and the conclusion drawn from this work can be generalized to other similar cases.
The construct threat could be the hyperparameter settings and the non-determinism of LLMs. We directly use the default settings suggested by Hugging Face and OpenAI’s API to build or access LLMs. Besides, we release the datasets and prompts used in our experiments to support reproducing results reported in our work and for future research.
@@ 8. Conclusion
In this work, we conducted the first empirical study to explore the potential of fault detection methods for LLMs. Based on the experiments on four tasks including NLP and code tasks, four LLMs, and nine fault detection methods, we found LLMs are overconfident with their predictions and simple methods such as Margin performs the best in revealing faults in LLMs. To enhance the detection performance, we proposed MuCS, which uses mutation analysis to smooth the prediction confidence of LLMs. Specifically, we utilized text transformation and code refactoring techniques to mutate the inputs and collect the averaged output probabilities of all mutants as the final prediction confidence of LLMs. Evaluation results demonstrated that MuCS significantly increased the diversity of output distribution of LLMs and enhanced the effectiveness of fault detection by up to 70.53% in terms of test relative coverage score.